Total number of protons and electrons present in 1g Co2

Grade:11

1 Answers

Khimraj
5 years ago
A atom has equal no of proton and electron.
proton in one atom of C = 6
proton in one atom of O = 8
proton in one molecule of CO2 = 6+ 8*2 = 22
1 gm CO2 = 1/44 mole
SO 
proton and electron in 1 mole of CO2 = (1/44)*22*6.022*1023
= 3.011*1023
Hope it clears.